Write and solve an equation to find the value of the variable.

5.2, 8.3, 8.5, 7.7, 7.8, y; mean = 7.1




1
Expert's answer
2022-03-05T07:07:34-0500

By definition of the mean, as there are 6 elements in total, we should have :

"\frac{5.2+8.3+8.5+7.7+7.8+y}{6}=7.1"

This gives us an equation for "y" which can be solved to find :

"5.2+8.3+8.5+7.7+7.8+y=6\times 7.1 \\ 
37.5+y=42.6 \\ 
y=42.6-37.5 \\ 
y=5.1"